OpenStudy (anonymous):

When you simplify each side of the equation 5n – 2(3 – n) = 4(2n – 3), you will get the equivalent equation 7n – 6 = 8n - 12. True False

OpenStudy (johnweldon1993):

lets see 5n - 2(3-n) = 4(2n - 3) 5n -6 + 2n = 8n - 12 7n - 6 = 8n - 12 true! :)

OpenStudy (anonymous):

@johnweldon1993 Can you help me with one more please? After isolating the absolute value in the equation -2|5n + 10| = -30, you would see that there is “no solution” for this problem. True False

OpenStudy (anonymous):

Oops, sorry I had to double click!

OpenStudy (johnweldon1993):

well lets check this one now -2|5n + 10| = -30 as it says you want to isolate the absolute value right? so how do we get it by itself? ***hint...what do we do with that 2 on the outside?

OpenStudy (johnweldon1993):

-2* sorry

OpenStudy (anonymous):

we add it to both sides?

OpenStudy (johnweldon1993):

notice that it is being multiplied to the absolute value.....so the opposite of multiply is?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

true so it make -2|5n+10|=-30 |5n+10|=15

OpenStudy (johnweldon1993):

correct....so we would divide both sides by -2

OpenStudy (johnweldon1993):

and as @nader1 has shown that will leave you with |5n + 10| = 15 this can be split into 2 things right? 5n + 10 = 15 and 5n + 10 = -15 so lets check it 5n + 10 = 15 -10 -10 5n = 5 n = 1 5n + 10 = -15 -10 -10 5n = -25 n = -5 there are solutions

OpenStudy (anonymous):

now you can simplify that into n=1 or n=-5

OpenStudy (anonymous):

yea true !

OpenStudy (anonymous):

Wow, that was a fantastic explanation @johnweldon1993! Thank you so much :) And thank you as well @nader1!

OpenStudy (johnweldon1993):

anytime :) you're very welcome!

OpenStudy (anonymous):

he is faster than me ^^ but anytime :)

OpenStudy (anonymous):

i at least enjoy it :)

OpenStudy (anonymous):

Haha, yea. Have a good night!

OpenStudy (johnweldon1993):

lol....I enjoy it as well :)

OpenStudy (anonymous):

hehe we are waiting for next time :D